A block of mass M is connected to another block of mass 2 by a massless spring force constant K the blocks are resting on a smooth horizontal plane the blocks are initially at rest and the spring and stitch then the constant force f is applied on block M1 to pull it then force on the block M2

Grade:12

1 Answers

Arun
5 years ago
Dear student
 
 

When 2 Blocks are in Contact -



F-f=m_{1}a

f=m_{2}a

- wherein

a=\frac{F}{m_{1}+m_{2}}

f=\frac{m_{2}F}{m_{1}+m_{2}}

 

 

 

: Acceleration \: of\: the\: system\: a= \frac{F}{m+M}

Force \: on \: block\: of \: mass\: m= ma= \frac{mF}{m+M}

 

In case of any query please feel free to ask again.

 

Regards

Arun (askIITians forum expert)